# Functional Central Limit Theorem for the simultaneous subgraph count of dynamic Erdős-Rényi random graphs

### v1

@@ 3. Graph notations
    Finally, for any graphs ${H},{H^{\star}}$ and $g\in{\tt CS}({H},{H^{\star}})$, denoting $\mathcal{N}=\mathcal{V}({H})+\mathcal{V}({H^{\star}})-\mathcal{V}(g)$, we define
    $\mathcal{C}({H},{H^{\star}},g):=\sum_{\begin{subarray}{c}g_{1}\in{\tt G}_{\mathcal{N}}({H})\\ g_{2}\in{\tt G}_{\mathcal{N}}({H^{\star}})\\ g_{1}\cap g_{2}\sim g\end{subarray}}\frac{1}{\mathcal{N}!}$ (3.14)
    Then, for any $N\geqslant\mathcal{N}$ we can see that the number of embedded graphs $H$ and ${H^{\star}}$ into $K_{N}$ such that their intersection is isomorphic to $g$ equals
    $\sum_{\begin{subarray}{c}g_{1}\in{\tt G}_{N}({H})\\ g_{2}\in{\tt G}_{N}({H^{\star}})\\ g_{1}\cap g_{2}\sim g\end{subarray}}1=\frac{N!}{(N-\mathcal{N})!}\mathcal{C}({H},{H}^{\star},g).$
@@
    Lemma 3.4.
    Let $H$, $H^{\star}$ be two graphs which satisfy (3.4). If $H\cong H^{\star}$ and $g\in{\tt OCS}(H,H^{\star})$, we have the following representation for the constant $\mathcal{C}(H,H^{\star},g)$:
@@
    where for any graph $H^{\prime}$ and any $g\in{\tt CS}(H^{\prime})$ the constant $\mathcal{S}(H^{\prime},g)$ is the total number of different subgraphs of $H^{\prime}$ which are isomorphic to $g$, i.e.,
    $\mathcal{S}(H^{\prime},g)=\sum_{g_{0}\,\sqsubset\,{H^{\star}}}\mathbb{I}\{g_{0}\sim g\},$ (3.15)
    where $H^{\star}\,\sqsubset\,K_{\mathcal{V}(H^{\prime})}$ is any embedded subgraph of $K_{\mathcal{V}(H^{\prime})}$ which is isomorphic to $H^{\prime}$.
@@
4. Main result
In this section we state our functional central limit theorem for a vector of subgraph count processes. To this end we fix, for a given $m\in{\mathbb{N}}$, (deterministic) graphs ${H}_{1},\ldots,{H}_{m}$, which satisfy Assumption 3.1; think for instance of $m=2$ and ${H}_{1}$ corresponding to edges and ${H}_{2}$ to triangles. Recall that from (3.4) we know that for any $i\in\langle m\rangle$ and any $g\subset{H}_{i}$ with at least one vertex
@@ 4. Main result
where, for $i\in\langle m\rangle$, $X_{N,i}(t)$ records the number of inclusions of the graph ${H}_{i}$ in $G_{N}(t)$. Using (3.2) the marginals of the process $X_{N}(t)$ can be alternatively defined via
$X_{N,i}(t)=\sum_{g\in{\tt G}_{N}({H}_{i})}\mathbb{I}\{g{\,\sqsubset\,}G_{N}(t)\}=\sum_{g\in{\tt G}_{N}({H}_{i})}\prod_{(u,v)\in{\tt E}(g)}a_{N,u,v}(t).$ (4.2)
For each graph $g$ we define a centered Gaussian process $(X^{\star}_{g}(t))_{t\in[0,T]}$, characterized via the covariance function
@@
in addition, all processes $(X^{\star}_{g}(t))_{t\in[0,T]}$ are independent (across the graphs $g$). We can now state our main result.
@@
Let $G_{N}(t)$ for $t\in[0,T]$ be a dynamic Erdős-Rényi random graph with $N$ vertices which satisfies Assumptions 2.1 and 2.4, and let $H_{1},\ldots,H_{m}$ be graphs which satisfy Assumption 3.1. Then, as $N\to\infty$, the process $\boldsymbol{X}_{N}(\cdot)$ defined in (4.2) satisfies the following limiting theorem:
@@
$\Sigma_{i,j}(s,t)=\mathbb{I}\{{H}_{i}\cong{H}_{j}\}\sum_{g\in{\tt OCS}({H}_{i},{H}_{j})}\frac{\mathcal{A}(g)\,\mathcal{S}({H}_{i},g)\,\mathcal{S}({H}_{j},g)}{\mathcal{A}({H}_{i})\,\mathcal{A}({H}_{j})}\frac{\bigl{(}p_{\star}(s)\bigr{)}^{\mathcal{E}({H}_{i})}\bigl{(}p_{\star}(t)\bigr{)}^{\mathcal{E}({H}_{j})}}{\bigl{(}p_{\star}(s)\,p_{\star}(t)\bigr{)}^{\mathcal{E}(g)}}\kappa^{\mathcal{E}(g)}_{\star}(s,t).$
[…]
@@
Let $G_{N}(t)$ for $t\in[0,T]$ be a dynamic Erdős-Rényi random graph with $N$ vertices, which satisfies Assumption 2.4 (for $\varrho_{N}=1$), and assume that the edge process $a_{N}(\cdot)$ does not depend on $N$. Let $H_{1},\ldots,H_{m}$ be graphs with at least one edge. Then, the process $\boldsymbol{X}_{N}(\cdot)$ defined in (4.2) satisfies the following limiting theorem: as $N\to\infty$,
$\left(\frac{X_{N,1}(t)-\mathbb{E}\left\{X_{N,1}(t)\right\}}{N^{\mathcal{V}(H_{1})-1}},\ldots,\frac{X_{N,m}(t)-\mathbb{E}\left\{X_{N,m}(t)\right\}}{N^{\mathcal{V}(H_{m})-1}}\right)^{\top}\to\boldsymbol{X}(t),$
@@
and where $p(t)=p_{N}(t)$, $\kappa(s,t)=\kappa_{N}(s,t)$ (which are independent of $N$ due to our assumption), and $X^{\star}(\cdot)$ is a centered Gaussian process characterized via its covariance function
@@
The next step is to describe the sets ${\tt OCS(H_{1})}$ and ${\tt OCS(H_{2})}$. The first one is easy to find: ${\tt OCS(H_{1})}=\{g_{1}\}$, where $g_{1}$ is an edge. However, for the second one we need to distinguish three cases:
@@
where $g_{2}$ is a triangle. It is readily verified that
@@
[…]
@@
$\mathbb{E}\left\{X_{N,i}(t)\right\}=\frac{N!\,p_{N}^{\mathcal{E}({H}_{i})}(t)}{(N-\mathcal{V}({H}_{i}))!\mathcal{A}({H}_{i})}$ (5.2)
@@
We postpone the proof of Lemma 5.1 to Section 6. To link the results of Lemma 5.1 with the covariance function of the process $\boldsymbol{X}(\cdot)$, we rely on the following proposition, proven in the appendix.
Proposition 5.3.
@@
$\mathbb{E}\left\{\left\lvert\boldsymbol{X}^{\star}_{N}(t)-\boldsymbol{X}^{\star}_{N}(s)\right\rvert^{2}\left\lvert\boldsymbol{X}^{\star}_{N}(s)-\boldsymbol{X}^{\star}_{N}(r)\right\rvert^{2}\right\}\leqslant\left\lvert F(t)-F(r)\right\rvert^{2}$ (5.4)
@@
The proof of this lemma being provided in Section 8. We obtain the claimed convergence by combining [3, Theorem 13.5] with Lemma 5.5, Lemma 5.4 and Proposition 5.3.
6. Proof of Lemma 5.1
@@
$\mathbb{E}\left\{X_{N,i}(t)\right\}$ $=\sum_{g\in{\tt G}_{N}({H_{i}})}\mathbb{E}\left\{\mathbb{I}\{g{\,\sqsubset\,}G_{N}(t)\}\right\}$
$=\sum_{g\in{\tt G}_{N}({H_{i}})}\mathbb{P}\left\{\prod_{(u,v)\in{\tt E}(g)}a_{u,v}(t)=1\right\}=\sum_{g\in{\tt G}_{N}({H_{i}})}\prod_{(u,v)\in{\tt E}(g)}\mathbb{P}\left\{a_{N,u,v}(t)=1\right\}$
@@
$\operatorname{Cov}\bigl{(}X_{N,i}(s),X_{N,j}(t)\bigr{)}$ $=\operatorname{Cov}\left(\sum_{g\in{\tt G}_{N}({H_{i}})}\mathbb{I}\{g{\,\sqsubset\,}G_{N}(s)\},\sum_{g\in{\tt G}_{N}({H_{j}})}\mathbb{I}\{g{\,\sqsubset\,}G_{N}(t)\}\right)$
$=\sum_{\begin{subarray}{c}g_{1}\in{\tt G}_{N}({H_{i}})\\ g_{2}\in{\tt G}_{N}({H_{j}})\end{subarray}}\operatorname{Cov}\bigl{(}\mathbb{I}\{g_{1}{\,\sqsubset\,}G_{N}(s)\},\mathbb{I}\{g_{2}{\,\sqsubset\,}G_{N}(t)\}\bigr{)}.$ (6.1)
Using (3.2) and the fact that the processes $a_{N,u,v}(\cdot)$ are independent for different pairs $(u,v)$, we obtain that the dependence of the random variables $\mathbb{I}\{g_{1}{\,\sqsubset\,}G_{N}(s)\}$ and $\mathbb{I}\{g_{2}{\,\sqsubset\,}G_{N}(t)\}$ can be expressed in terms of the intersection ${\tt E}(g_{1})\cap{\tt E}(g_{2})={\tt E}(g_{1}\cap g_{2})$, i.e.,
$\operatorname{Cov}\bigl{(}\mathbb{I}\{g_{1}{\,\sqsubset\,}G_{N}(t)\},$ $\mathbb{I}\{g_{2}{\,\sqsubset\,}G_{N}(s)\}\bigr{)}=\operatorname{Cov}\left(\prod_{(u,v)\in{\tt E}(g_{1})}a_{N,u,v}(t),\prod_{(u,v)\in{\tt E}(g_{2})}a_{N,u,v}(s)\right)$
@@
×$\qquad\times\operatorname{Cov}\left(\prod_{(u,v)\in{\tt E}(g_{1})\cap{\tt E}(g_{2})}a_{N,u,v}(t),\prod_{(u,v)\in{\tt E}(g_{1})\cap{\tt E}(g_{2})}a_{N,u,v}(s)\right)$
@@
×$\qquad\times\left(\bigl{(}\kappa_{N}(t,s)+p_{N}(t)p_{N}(s)\bigr{)}^{\mathcal{E}(g_{1}\cap g_{2})}-\bigl{(}p_{N}(t)p_{N}(s)\bigr{)}^{\mathcal{E}(g_{1}\cap g_{2})}\right).$ (6.2)
Hence, each term in the sum (6.1) depends on the ‘structure’ of the intersection $g_{1}\cap g_{2}$. We show that the dominant part of the asymptotics of (6.1) is due to such pairs $g_{1},g_{2}$ that are such that their intersection is isomorphic to the optimal common subgraphs of $H_{i}$ and $H_{j}$.
@@
$\operatorname{Cov}$ $\bigl{(}\mathbb{I}\{g_{1}{\,\sqsubset\,}G_{N}(s)\},\mathbb{I}\{g_{2}{\,\sqsubset\,}G_{N}(t)\}\bigr{)}=$
@@
$\frac{\operatorname{Cov}\bigl{(}X_{N,i}(t),X_{N,j}(s)\bigr{)}}{\varrho_{N}^{\mathcal{E}({H}_{i})+\mathcal{E}({H}_{j})}\bigl{(}p_{\star}(s)\bigr{)}^{\mathcal{E}({H}_{i})}\bigl{(}p_{\star}(t)\bigr{)}^{\mathcal{E}({H}_{j})}}$ $=(1+o(1))\sum_{\begin{subarray}{c}g_{\cap}\in{\tt CS}({H}_{i},{H}_{j}),\\ \mathcal{V}{(g_{\cap})\geqslant 1}\end{subarray}}\sum_{\begin{subarray}{c}g_{1}\in{\tt G}_{N}({H}_{i})\\ g_{2}\in{\tt G}_{N}({H}_{j})\\ g_{1}\cap g_{2}\sim{g_{\cap}}\end{subarray}}\varrho_{N}^{-\mathcal{E}(g_{\cap})}\left(\frac{\kappa_{\star}(s,t)}{p_{\star}(t)p_{\star}(s)}\right)^{\mathcal{E}(g_{\cap})}$
$=(1+o(1))\sum_{\begin{subarray}{c}g_{\cap}\in{\tt CS}({H}_{i},{H}_{j}),\\ \mathcal{V}(g_{\cap}\geqslant 1)\end{subarray}}\mathcal{C}({H}_{i},{H}_{j},g_{\cap})N^{\mathcal{V}({H}_{i})+\mathcal{V}({H}_{j})-\mathcal{V}(g_{\cap})}\varrho_{N}^{-\mathcal{E}(g_{\cap})}$
×$\qquad\qquad\qquad\qquad\qquad\times\left(\frac{\kappa_{\star}(s,t)}{p_{\star}(t)p_{\star}(s)}\right)^{\mathcal{E}(g_{\cap})}$
@@
×$\qquad\times\sum_{\begin{subarray}{c}g_{\cap}\in{\tt CS}({H}_{i},{H}_{j}),\\ \mathcal{V}(g_{\cap}\geqslant 1)\end{subarray}}\mathcal{C}({H}_{i},{H}_{j},g_{\cap})\frac{\mathcal{F}_{N}(g_{\cap})}{\mathcal{F}_{N}^{\rm opt}({H}_{i},{H}_{j})}\left(\frac{\kappa_{\star}(s,t)}{p_{\star}(t)p_{\star}(s)}\right)^{\mathcal{E}(g_{\cap})}$
@@
×$\qquad\qquad\qquad\times\sum_{\begin{subarray}{c}g_{\cap}\in{\tt OCS}({H}_{i},{H}_{j}),\\ \mathcal{V}(g_{\cap}\geqslant 1)\end{subarray}}\mathcal{C}({H}_{i},{H}_{j},g_{\cap})\left(\frac{\kappa_{\star}(s,t)}{p_{\star}(t)p_{\star}(s)}\right)^{\mathcal{E}(g_{\cap})},$
@@
where $\mathfrak{e}$ is an edge. We observe that the covariance representation obtained in (6.2) depends only on the number of common edges in the graphs $g_{1}$ and $g_{2}$. Hence we can classify all the pairs $g_{1},g_{2}$ on the right hand side of (6.1) by $k=\mathcal{E}(g_{1}\cap g_{2})$. We thus obtain, as $N\to\infty$,
$\operatorname{Cov}\bigl{(}X_{N}(t),X_{N}(s)\bigr{)}$ $=\sum_{k=0}^{\min(\mathcal{E}({H}_{i}),\mathcal{E}({H}_{j}))}\sum_{\begin{subarray}{c}g_{1}\in{\tt G}_{N}({H}_{i})\\ g_{2}\in{\tt G}_{N}({H}_{j})\\ \mathcal{E}(g_{1}\cap g_{2})=k\end{subarray}}\bigl{(}p_{N}(t)\bigr{)}^{\mathcal{E}({H}_{i})-k}\bigl{(}p_{N}(s)\bigr{)}^{\mathcal{E}({H}_{j})-k}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times\left(\bigl{(}\kappa_{N}(t,s)+p_{N}(t)p_{N}(s)\bigr{)}^{k}-\bigl{(}p_{N}(t)p_{N}(s)\bigr{)}^{k}\right)$
@@
×$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times\left(\bigl{(}\kappa_{\star}(t,s)+p_{\star}(t)p_{\star}(s)\bigr{)}^{k}-\bigl{(}p_{\star}(t)p_{\star}(s)\bigr{)}^{k}\right),$
@@
Hence, applying this asymptotic expression, we obtain that as $N\to\infty$
$\operatorname{Cov}\bigl{(}X_{N}(t),X_{N}(s)\bigr{)}$ $=(1+o(1))C({H}_{i},{H}_{j},\mathfrak{e})N^{\mathcal{V}({H}_{i})+\mathcal{V}({H}_{j})-2}\kappa_{\star}(t,s)\bigl{(}p_{\star}(t)\bigr{)}^{\mathcal{E}({H}_{i})-1}\bigl{(}p_{\star}(s)\bigr{)}^{\mathcal{E}({H}_{j})-1}.$ (6.4)
Hence, claim (5.3) follows as well in the case that $\varrho_{N}=1$.

### v2

@@ 3 Graph notations
    Finally, for any graphs ${H},{H^{\star}}{\in{\tt G}_{\infty}}$ and $g\in{\tt CS}({H},{H^{\star}})$, denoting ${\mathcal{N}(H,H^{\star},g)}=\mathcal{V}({H})+\mathcal{V}({H^{\star}})-\mathcal{V}(g)$, we define
    $\mathcal{C}({H},{H^{\star}},g):=\sum_{\begin{subarray}{c}g_{1}\in{\tt G}_{{\mathcal{N}(H,H^{\star},g)}}({H})\\ g_{2}\in{\tt G}_{{\mathcal{N}(H,H^{\star},g)}}({H^{\star}})\\ {\pi(g_{1}\cap g_{2})=g}\end{subarray}}\frac{1}{{\mathcal{N}(H,H^{\star},g)}!}{.}$ (3.13)
    Then, for any $N\geqslant{\mathcal{N}(H,H^{\star},g)}$ we can see that the number of pairs of graphs $g_{1},g_{2}\in{\tt G}_{\infty}$ such that $g_{1}\sim H,\,g_{2}\sim{H^{\star}}$, ${\tt V}(g_{1}),{\tt V}(g_{2})\subset\langle N\rangle$ and $\pi(g_{1}\cap g_{2})=g$ equals
    $\sum_{\begin{subarray}{c}g_{1}\in{\tt G}_{N}({H})\\ g_{2}\in{\tt G}_{N}({H^{\star}})\\ {\pi(g_{1}\cap g_{2})\sim g}\end{subarray}}1=\frac{N!}{(N-{\mathcal{N}(H,H^{\star},g)})!}\mathcal{C}({H},{H}^{\star},g).$
@@
    Lemma 2.
    Let $H,\,H^{\star}{\in{\tt G}_{\infty}}$ be two graphs which satisfy Assumption 3.1. If $H\cong H^{\star}$ and $g\in{\tt OCS}(H,H^{\star})$, we have the following representation for the constant $\mathcal{C}(H,H^{\star},g)$:
@@
    where the constants $\mathcal{S}(H,g)$ and $\mathcal{S}(H^{\star},g)$ are defined in (3.3).
@@
4 Main result
In this section we state our functional central limit theorem for a vector of subgraph count processes. To this end, we fix, for a given $m\in{\mathbb{N}}$, (deterministic) graphs ${H}_{1},\ldots,{H}_{m}{\in{\tt G}_{\infty}}$, which satisfy Assumption 3.1; think for instance of $m=2$ and ${H}_{1}$ corresponding to edges and ${H}_{2}$ to triangles. Recall that from (3.5) we know that for any $i\in\langle m\rangle$ and any $g\subset{H}_{i}$ with at least one vertex
@@ 4 Main result
where, for $i\in\langle m\rangle$, $X_{N,i}(t)$ records the number of subgraphs of $G_{N}(t)$ which are isomorphic to ${H}_{i}$. Using (3.2) the marginals of the process ${\boldsymbol{X}}_{N}(t)$ can be alternatively defined via
$X_{N,i}(t)=\sum_{g\in{\tt G}_{N}({H}_{i})}\mathbb{I}\{g{\,{\subset}\,}G_{N}(t)\}=\sum_{g\in{\tt G}_{N}({H}_{i})}\prod_{(u,v)\in{\tt E}(g)}a_{N,u,v}(t).$ (4.2)
For each graph pattern $g\in{\tt G}_{\sim}$ we define a centered Gaussian process $(X^{\star}_{g}(t))_{t\in[0,T]}$, characterized via the covariance function
@@
in addition, all processes $(X^{\star}_{g}(t))_{t\in[0,T]}$ are independent (across the graph patterns $g$). We can now state our main result.
@@
Let $G_{N}(t)$ for $t\in[0,T]$ be a dynamic Erdős–Rényi random graph with $N$ vertices which satisfies Assumptions 2.1 and 2.2, and let $H_{1},\ldots,H_{m}{\in{\tt G}_{\infty}}$ be graphs which satisfy Assumption 3.1. Then, as $N\to\infty$, the process $\boldsymbol{X}_{N}(\cdot)$ defined in (4.2) satisfies the following limiting theorem:
@@
$\Sigma_{i,j}(s,t)=\mathbb{I}\{{H}_{i}\cong{H}_{j}\}\sum_{g\in{\tt OCS}({H}_{i},{H}_{j})}\frac{\mathcal{A}(g)\,\mathcal{S}({H}_{i},g)\,\mathcal{S}({H}_{j},g)}{\mathcal{A}({H}_{i})\,\mathcal{A}({H}_{j})}\frac{\bigl(p_{\star}(s)\bigr)^{\mathcal{E}({H}_{i})}\bigl(p_{\star}(t)\bigr)^{\mathcal{E}({H}_{j})}}{\bigl(p_{\star}(s)\,p_{\star}(t)\bigr)^{\mathcal{E}(g)}}\kappa^{\mathcal{E}(g)}_{\star}(s,t)$
for $s,t\geqslant 0$. $\Diamond$}
@@
Let $G_{N}(t)$ for $t\in[0,T]$ be a dynamic Erdős–Rényi random graph with $N$ vertices, which satisfies Assumption 2.2 (for $\varrho_{N}=1$), and assume that the edge process $a_{N}(\cdot)$ does not depend on $N$. Let $H_{1},\ldots,H_{m}\in{\tt G}_{\infty}$ be graphs with at least one edge. Then, the process $\boldsymbol{X}_{N}(\cdot)$ defined in (4.2) satisfies the following limiting theorem: as $N\to\infty$,
$\left(\frac{X_{N,1}(t)-\mathbb{E}\left\{X_{N,1}(t)\right\}}{N^{\mathcal{V}(H_{1})-1}},\ldots,\frac{X_{N,m}(t)-\mathbb{E}\left\{X_{N,m}(t)\right\}}{N^{\mathcal{V}(H_{m})-1}}\right)^{\top}\to\boldsymbol{X}(t)$
@@
and where $p(t)=p_{N}(t)$, $\kappa(s,t)=\kappa_{N}(s,t)$ (which are independent of NN due to our assumption), and $X^{\star}(\cdot)$ is a centered Gaussian process characterized via its covariance function
@@
The next step is to describe the sets $\tt OCS(H_{1})$} and $\tt OCS(H_{2})$}. The first one is easy to find: ${\tt OCS(H_{1})}=\{g_{1}\}$, where $g_{1}\in{\tt G}_{\sim}$} is the pattern of an edge. However, for the second one we need to distinguish three cases:
@@
where $g_{2}\in{\tt G}_{\sim}$} is the pattern of a triangle. It is readily verified that
@@
$\left(\frac{X_{N,1}(t)-\mathbb{E}\left\{X_{N,1}(t)\right\}}{N^{2-3\alpha/2}},\frac{X_{N,2}(t)-\mathbb{E}\left\{X_{N,2}(t)\right\}}{N^{\max(2-5\alpha/2,\,3/2-3\alpha/2)}}\right)^{\top}\to(X_{1}(t),X_{2}(t))^{\top},$
@@
$\mathbb{E}\left\{X_{N,i}(t)\right\}=\frac{N!\,p_{N}^{\mathcal{E}({H}_{i})}(t)}{(N-\mathcal{V}({H}_{i}))!\mathcal{A}({H}_{i})}{,}$ (5.2)
@@
We postpone the proof of Lemma 3 to Section 6. To link the results of Lemma 3 with the covariance function of the process $\boldsymbol{X}(\cdot)$, we rely on the following proposition, proven in the appendix.
Proposition 4.
@@
$\mathbb{E}\left\{\left\lvert\boldsymbol{X}^{\star}_{N}(t)-\boldsymbol{X}^{\star}_{N}(s)\right\rvert^{2}\left\lvert\boldsymbol{X}^{\star}_{N}(s)-\boldsymbol{X}^{\star}_{N}(r)\right\rvert^{2}\right\}\leqslant\left\lvert F(t)-F(r)\right\rvert^{2}{,}$ (5.4)
@@
The proof of this lemma being provided in Section 8. We obtain the claimed convergence by combining [3, Theorem 13.5] with Lemmas 5 and 6 and Proposition 4.
6 Proof of Lemma 3
@@
$\mathbb{E}\left\{X_{N,i}(t)\right\}$ $=\sum_{g\in{\tt G}_{N}({H_{i}})}\mathbb{E}\left\{\mathbb{I}\{g{\,{\subset}\,}G_{N}(t)\}\right\}$
$=\sum_{g\in{\tt G}_{N}({H_{i}})}\mathbb{P}\left\{\prod_{(u,v)\in{\tt E}(g)}a_{{N,u,v}}(t)=1\right\}=\sum_{g\in{\tt G}_{N}({H_{i}})}\prod_{(u,v)\in{\tt E}(g)}\mathbb{P}\left\{a_{N,u,v}(t)=1\right\}$
@@
$\operatorname{Cov}\bigl(X_{N,i}(s),X_{N,j}(t)\bigr)$ $=\operatorname{Cov}\left(\sum_{g\in{\tt G}_{N}({H_{i}})}\mathbb{I}\{g{\,{\subset}\,}G_{N}(s)\},\sum_{g\in{\tt G}_{N}({H_{j}})}\mathbb{I}\{g{\,{\subset}\,}G_{N}(t)\}\right)$
$=\sum_{\begin{subarray}{c}g_{1}\in{\tt G}_{N}({H_{i}})\\ g_{2}\in{\tt G}_{N}({H_{j}})\end{subarray}}\operatorname{Cov}\bigl(\mathbb{I}\{g_{1}{\,{\subset}\,}G_{N}(s)\},\mathbb{I}\{g_{2}{\,{\subset}\,}G_{N}(t)\}\bigr).$ (6.1)
Using (3.2) and the fact that the processes $a_{N,u,v}(\cdot)$ are independent for different pairs $(u,v)$, we obtain that the dependence of the random variables $\mathbb{I}\{g_{1}{\,{\subset}\,}G_{N}(s)\}$ and $\mathbb{I}\{g_{2}{\,{\subset}\,}G_{N}(t)\}$ can be expressed in terms of the intersection ${\tt E}(g_{1})\cap{\tt E}(g_{2})={\tt E}(g_{1}\cap g_{2})$, i.e.,
$\operatorname{Cov}\bigl(\mathbb{I}\{g_{1}{\,{\subset}\,}G_{N}(t)\},$ $\mathbb{I}\{g_{2}{\,{\subset}\,}G_{N}(s)\}\bigr)=\operatorname{Cov}\left(\prod_{(u,v)\in{\tt E}(g_{1})}a_{N,u,v}(t),\prod_{(u,v)\in{\tt E}(g_{2})}a_{N,u,v}(s)\right)$
@@
$\qquad\times\operatorname{Cov}\left(\prod_{(u,v)\in{\tt E}(g_{1})\cap{\tt E}(g_{2})}a_{N,u,v}(t),\prod_{(u,v)\in{\tt E}(g_{1})\cap{\tt E}(g_{2})}a_{N,u,v}(s)\right)$
@@
$\qquad\times\left(\bigl(\kappa_{N}(t,s)+p_{N}(t)p_{N}(s)\bigr)^{\mathcal{E}(g_{1}\cap g_{2})}-\bigl(p_{N}(t)p_{N}(s)\bigr)^{\mathcal{E}(g_{1}\cap g_{2})}\right).$ (6.2)
Hence, each term in the sum (6.1) depends on the pattern of the intersection $g_{1}\cap g_{2}$. We show that the dominant part of the asymptotics of (6.1) is due to such pairs $g_{1},g_{2}$ that are such that their intersection is isomorphic to the optimal common subgraph pattern of $H_{i}$ and $H_{j}$.
@@
$\operatorname{Cov}$ $\bigl(\mathbb{I}\{g_{1}{\,{\subset}\,}G_{N}(s)\},\mathbb{I}\{g_{2}{\,{\subset}\,}G_{N}(t)\}\bigr)=$
@@
$\frac{\operatorname{Cov}\bigl(X_{N,i}(t),X_{N,j}(s)\bigr)}{\varrho_{N}^{\mathcal{E}({H}_{i})+\mathcal{E}({H}_{j})}\bigl(p_{\star}(s)\bigr)^{\mathcal{E}({H}_{i})}\bigl(p_{\star}(t)\bigr)^{\mathcal{E}({H}_{j})}}$ $=(1+o(1))\sum_{\begin{subarray}{c}g_{\cap}\in{\tt CS}({H}_{i},{H}_{j}),\\ \mathcal{V}{(g_{\cap})\geqslant 1}\end{subarray}}\sum_{\begin{subarray}{c}g_{1}\in{\tt G}_{N}({H}_{i})\\ g_{2}\in{\tt G}_{N}({H}_{j})\\ {\pi(g_{1}\cap g_{2})={g_{\cap}}}\end{subarray}}\varrho_{N}^{-\mathcal{E}(g_{\cap})}\left(\frac{\kappa_{\star}(s,t)}{p_{\star}(t)p_{\star}(s)}\right)^{\mathcal{E}(g_{\cap})}$
$=(1+o(1))\sum_{\begin{subarray}{c}g_{\cap}\in{\tt CS}({H}_{i},{H}_{j}),\\ \mathcal{V}(g_{\cap})\geqslant 1\end{subarray}}\mathcal{C}({H}_{i},{H}_{j},g_{\cap})N^{\mathcal{V}({H}_{i})+\mathcal{V}({H}_{j})-\mathcal{V}(g_{\cap})}\varrho_{N}^{-\mathcal{E}(g_{\cap})}$
$\qquad\qquad\qquad\qquad\qquad\times\left(\frac{\kappa_{\star}(s,t)}{p_{\star}(t)p_{\star}(s)}\right)^{\mathcal{E}(g_{\cap})}$
@@
$\qquad\times\sum_{\begin{subarray}{c}g_{\cap}\in{\tt CS}({H}_{i},{H}_{j}),\\ \mathcal{V}(g_{\cap})\geqslant 1\end{subarray}}\mathcal{C}({H}_{i},{H}_{j},g_{\cap})\frac{\mathcal{F}_{N}(g_{\cap})}{\mathcal{F}_{N}^{\rm opt}({H}_{i},{H}_{j})}\left(\frac{\kappa_{\star}(s,t)}{p_{\star}(t)p_{\star}(s)}\right)^{\mathcal{E}(g_{\cap})}$
@@
$\qquad\qquad\qquad\times\sum_{\begin{subarray}{c}g_{\cap}\in{\tt OCS}({H}_{i},{H}_{j}),\\ \mathcal{V}(g_{\cap})\geqslant 1\end{subarray}}\mathcal{C}({H}_{i},{H}_{j},g_{\cap})\left(\frac{\kappa_{\star}(s,t)}{p_{\star}(t)p_{\star}(s)}\right)^{\mathcal{E}(g_{\cap})},$
@@
where $\mathfrak{e}$ is the pattern of an edge. We observe that the covariance representation obtained in (6.2) depends only on the number of common edges in the graphs $g_{1}$ and $g_{2}$. Hence we can classify all the pairs $g_{1},g_{2}$ on the right hand side of (6.1) by $k=\mathcal{E}(g_{1}\cap g_{2})$. We thus obtain, as $N\to\infty$,
$\operatorname{Cov}\bigl({X_{N,i}(t),X_{N,i}(s)}\bigr)$ $=\sum_{k=0}^{\min(\mathcal{E}({H}_{i}),\mathcal{E}({H}_{j}))}\sum_{\begin{subarray}{c}g_{1}\in{\tt G}_{N}({H}_{i})\\ g_{2}\in{\tt G}_{N}({H}_{j})\\ \mathcal{E}(g_{1}\cap g_{2})=k\end{subarray}}\bigl(p_{N}(t)\bigr)^{\mathcal{E}({H}_{i})-k}\bigl(p_{N}(s)\bigr)^{\mathcal{E}({H}_{j})-k}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times\left(\bigl(\kappa_{N}(t,s)+p_{N}(t)p_{N}(s)\bigr)^{k}-\bigl(p_{N}(t)p_{N}(s)\bigr)^{k}\right)$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\times\left(\bigl(\kappa_{\star}(t,s)+p_{\star}(t)p_{\star}(s)\bigr)^{k}-\bigl(p_{\star}(t)p_{\star}(s)\bigr)^{k}\right),$
@@
Hence, applying this asymptotic expression, we obtain that, as $N\to\infty$,
$\operatorname{Cov}\bigl(X_{N}(t),X_{N}(s)\bigr)$ (6.4)
$\qquad=(1+o(1))C({H}_{i},{H}_{j},\mathfrak{e})N^{\mathcal{V}({H}_{i})+\mathcal{V}({H}_{j})-2}\kappa_{\star}(t,s)\bigl(p_{\star}(t)\bigr)^{\mathcal{E}({H}_{i})-1}\bigl(p_{\star}(s)\bigr)^{\mathcal{E}({H}_{j})-1}.$ (6.5)
We conclude that claim (5.3) follows as well in the case that $\varrho_{N}=1$.